OpenStudy (xbeenchantedx):

Using the Fundamental Counting Principle, how do i solve this word problem: your state issues license plates consisting of letters and numbers. there are 26 letters and 10 numbers. how many possible license plates can be issued with 2 letters followed by 3 numbers if letters and numbers must be unique?

OpenStudy (anonymous):

counting principle says multiply all the choices together you get \(26\times 26\times 10\times 10\times 10\)

OpenStudy (anonymous):

more succinctly written as \[26^2\times 10^3\]